# Average entropy of Gaussian mixtures

### v1

@@ Acknowledgements
We have $\mathbf{u}_{q}=\mathbf{r}/\sqrt{m_{2}}$, $\mathbf{u}_{q-1}=\mathbf{s}/\sqrt{m_{1}}$. This normalizes the variance of ${\mathbf{r}},{\mathbf{s}}$ to 1. Hence each component $r_{1},\ldots,r_{n}$, $s_{1},\ldots,s_{n}$ is an independent normal-distributed variable. We show that it is consistent to write ${\binom{{\mathbf{r}}}{{\mathbf{s}}}}=W{\binom{\tilde{\mathbf{r}}}{\tilde{\mathbf{s}}}}$, with $W$ as given in (34) and all the components $\tilde{r}_{1},\ldots,\tilde{r}_{n}$, $\tilde{s}_{1},\ldots,\tilde{s}_{n}$ independent and normal distributed. The change of variables can be written component-wise as $r_{\alpha}=W_{11}\tilde{r}_{\alpha}+W_{12}\tilde{s}_{\alpha}$, $s_{\alpha}=W_{21}\tilde{r}_{\alpha}+W_{22}\tilde{s}_{\alpha}$, for $\alpha\in\{1,\ldots,n\}$. A linear combination of normal-distributed variables has Gaussian distribution. Furthermore we have first-order statistics ${\mathbb{E}}[r_{\alpha}]=0$, ${\mathbb{E}}[s_{\alpha}]=0$. The second order statistics is given by:
@@
and ${\mathbb{E}}[r_{\alpha}r_{\beta}]=\delta_{\alpha\beta}$, ${\mathbb{E}}[s_{\alpha}s_{\beta}]=\delta_{\alpha\beta}$. The integration measure changes as ${\rm d}r_{\alpha}{\rm d}s_{\alpha}=|\det W|{\rm d}\tilde{r}_{\alpha}{\rm d}\tilde{s}_{\alpha}$, with
@@
$-6(q-1)[(q-2)ab^{2}B-2q(q-2)ab^{2}aa+q(q-2)ab^{2}bb+4q^{2}ab^{2}].$
Furthermore, the expression for $c_{1/2}$ and $c_{3/2}$ do not contribute to $S$ since they vanish under the expectation. We now compute analytically all the relevant terms needed.
@@
Direct substitution yields:
@@
From equation (149) and Theorem 2.7 we have:

### v2

@@ Acknowledgements
We have $\mathbf{u}_{q}=\mathbf{r}/\sqrt{m_{2}}$, $\mathbf{u}_{q-1}=\mathbf{s}/\sqrt{m_{1}}$. This normalizes the variance of ${\mathbf{r}},{\mathbf{s}}$ to 1. Hence, each component $r_{1},\ldots,r_{n}$, $s_{1},\ldots,s_{n}$ is an independent normal-distributed variable. We show that it is consistent to write ${\binom{{\mathbf{r}}}{{\mathbf{s}}}}=W{\binom{\tilde{\mathbf{r}}}{\tilde{\mathbf{s}}}}$, with $W$ as given in (30), and all the components $\tilde{r}_{1},\ldots,\tilde{r}_{n}$, $\tilde{s}_{1},\ldots,\tilde{s}_{n}$ are independent and normally distributed. The change of variables can be written component-wise as $r_{\alpha}=W_{11}\tilde{r}_{\alpha}+W_{12}\tilde{s}_{\alpha}$, $s_{\alpha}=W_{21}\tilde{r}_{\alpha}+W_{22}\tilde{s}_{\alpha}$, for $\alpha\in\{1,\ldots,n\}$. A linear combination of normal-distributed variables has a Gaussian distribution. Furthermore, we have first-order statistics ${\mathbb{E}}[r_{\alpha}]=0$, ${\mathbb{E}}[s_{\alpha}]=0$. The second-order statistics are given by
@@
and ${\mathbb{E}}[r_{\alpha}r_{\beta}]=\delta_{\alpha\beta}$, ${\mathbb{E}}[s_{\alpha}s_{\beta}]=\delta_{\alpha\beta}$. The integration measure changes as
${\rm d}r_{\alpha}{\rm d}s_{\alpha}=|\det W|{\rm d}\tilde{r}_{\alpha}{\rm d}\tilde{s}_{\alpha},$ (108)
with
@@
$-6(q-1)[(q-2)ab^{2}B-2q(q-2)ab^{2}aa+q(q-2)ab^{2}bb+4q^{2}ab^{2}].$ (118)
Furthermore, the expression for $c_{1/2}$ and $c_{3/2}$ do not contribute to $S$ since they vanish under the expectation. We now analytically compute all the relevant terms needed.
@@
Direct substitution yields
@@
From Equation (142) and Theorem 2.7, we have